OpenStudy (anonymous):

convert to degree measure -23/6 Pi * 180/ Pi

OpenStudy (lynfran):

is this the question \[\frac{ -23 }{ 6\pi }*\frac{ 180 }{ \pi }\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but the pie is next to the fraction

OpenStudy (anonymous):

-2 3/6 next to PI

OpenStudy (anonymous):

\[\frac{ -23 }{6 } \pi \times \frac{ 180 }{ \pi }\]

OpenStudy (lynfran):

ok 180/6= and we see that the pi would cancel out